# Models in Long-Context Interaction

### v1

@@ 3.4 Emotion QA
Dataset Construction.
Figure 4: The pipeline of Collaborative Emotional Modeling (CoEM). CoEM consists of five stages: Chunking, Initial Ranking, Multi-Agent Enrichment, Re-Ranking, and Emotional Ensemble Generation.
The annotation pipeline can be referred to in Figure 3. The construction process involves: i) expert-written questions targeting emotional understanding, ii) refinement of reference answers for clarity and consistency with F1-based evaluation, and iii) filtering based on model performance to exclude overly ambiguous or trivial examples. Through this series of manual annotation and selection process, we finally obtain 120 high-quality pairs of psychological knowledge questions and answers.
3.5 Emotion Summary
Task Design.
In this task, the model is required to summarize the following aspects from long-context psychological pathology reports: (i) causes, (ii) symptoms, (iii) treatment process, (iv) illness characteristics, and (v) treatment effects. After generating the model’s response, we employ GPT-4o to evaluate its factual consistency, completeness, and clarity with respect to the reference answer. These three evaluation criteria are validated in CPsyExam (Zhao et al. 2024).
Dataset Construction.
Drawing on CPsyCounR dataset, we first expand the experience and reflection section of the dataset to meet our requirements for long-context inputs. Next, psychology annotators label each sample across five standardized dimensions: i) Causes, ii) Symptoms, iii) Treatment Process, iv) Illness Characteristics, and v) Treatment Effect. Finally, by filtering samples based on format, content richness, and precision, we select a final set of 150 samples.
3.6 Emotion Expression
Task Design.
Dataset Construction.
We utilize the situations in the EmotionBench (Huang et al. 2024) to provide the model with a specific emotional context. Each instance specifies: i) an emotional category (e.g., anger), and ii) a psychologically meaningful trigger (e.g., being ignored in a group setting).
Method Model EC ED QA MC-4 ES $E$
Base GPT-4o-mini 28.50 16.42 48.61 3.75 4.14 86.77
GPT-4o 51.17 19.12 50.12 3.77 4.19 81.03
DeepSeek-V3 44.00 24.51 45.53 3.99 4.28 81.75
Qwen3-8B 38.50 18.14 44.75 3.97 4.21 73.40
Llama3.1-8B-Instruct 26.17 9.80 45.74 4.00 3.98 75.61
(Extended Comparison Models)
GPT-5 64.50 22.79 43.22 4.67 4.37 86.77
RAG GPT-4o-mini 38.33 21.57 50.72 3.78 4.19 80.41
GPT-4o 54.67 22.55 51.81 3.80 4.13 79.49
DeepSeek-V3 52.17 23.53 50.44 4.34 4.28 81.83
Qwen3-8B 39.67 19.12 44.34 4.14 4.20 73.28
Llama3.1-8B-Instruct 28.00 11.27 47.04 3.94 3.71 75.16
CoEM GPT-4o-mini 48.00 20.59 47.51 3.77 3.91 80.38
GPT-4o 52.83 25.00 47.24 3.81 4.02 80.41
DeepSeek-V3 54.33 23.04 46.52 4.34 4.12 82.83
Qwen3-8B 52.83 18.14 46.31 4.14 4.09 73.59
Llama3.1-8B-Instruct 38.17 11.27 44.79 4.00 3.60 75.71
Table 3: Experiment result across different prompting settings (Base, RAG, CoEM). EC represents Emotion Classification, ED represents Emotion Detection, QA represents Emotion QA, MC-4 represents the fourth stage of Emotion Conversation, ES represents Emotion Summary, and $E$ represents Emotion Expression.
(a) Impact of different CoEM-Sage models on MC-4.
(b) Impact of different CoEM-Sage models on ES.
Figure 5: Ablation experiments on CoEM-Sage models.
@@ 4 Collaborative Emotional Modeling
Figure 4 illustrates the pipeline of CoEM. To address EI tasks involving long contexts, we propose a hybrid retrieval-generation architecture that combines Retrieval-Augmented Generation (RAG) with modular multi-agent collaboration. The framework consists of five key stages:
@@
The input context is segmented into semantically coherent or token-length-constrained chunks. This enables efficient retrieval and minimizes irrelevant content during similarity estimation.
@@
A retrieval agent, implemented as CoEM-Rank, evaluates the relevance of each chunk to the query using dense semantic similarity. Top-ranked chunks are passed forward for enhancement.
@@
A reasoning agent called CoEM-Sage, functioning as a knowledge assistant, enriches the selected chunks by incorporating external knowledge or latent emotional signals. These signals, derived from psychological theories or curated priors, enhance emotional reasoning without introducing task-specific leakage.
@@
The enriched chunks are then re-evaluated by CoEM-Rank for both semantic relevance and emotional alignment. This ensures that the final input is both factually grounded and affectively coherent.
@@
The selected and enriched content, along with the prompt, is fed into a generation model denoted as CoEM-Core. This model (e.g., a long-context LLM or an instruction-tuned model) produces the final task-specific output, whether it be classification, summarization, or dialogue generation.
This modular approach encourages interpretability, emotional awareness, and task robustness. The CoEM setting encompasses all five stages, while the RAG setting only comprises Chunking, Re-Ranking, and Emotional Ensemble Generation. For the parameter settings and application details of RAG and CoEM, please refer to Appendix E.
@@ 5.1 Experiment Setup
In our experiments, for closed-source models, we choose GPT-4o-mini (OpenAI 2024a) and GPT-4o, while for open-source models, we select DeepSeek-V3 (DeepSeek-AI 2024), Llama3.1-8B-Instruct (Grattafiori et al. 2024), and Qwen3-8B (Team 2025). For tasks employing automatic evaluation, we adopt GPT-4o as the evaluator. Under the base setting, we compare a broader range of advanced open-source and closed-source models. We have currently evaluated the performance of GPT-5 (OpenAI 2025) and plan to include additional models in future experiments.
To accelerate inference, we use vllm library (Kwon et al. 2023) as the inference engine and set temperature=0.8 and top_p=0.9 for all models. For Qwen3-8B, we enable its thinking capabilities and manually remove the reasoning process between <think> and </think> to keep the answers concise. All our experiments are conducted on single A800 80G GPUs.
Model Stage 1 Stage 2 Stage 3 Stage 4 Avg
1-1 1-2 1-3 2-1 2-2 2-3 3-1 3-2 3-3 4-1 4-2 4-3
GPT-4o-mini 4.23 4.17 4.19 3.53 3.48 3.44 3.41 3.54 3.59 3.61 3.76 3.87 3.73
GPT-4o 4.06 4.17 4.19 3.47 3.44 3.41 3.44 3.54 3.60 3.63 3.75 3.92 3.72
Deepseek-V3 4.38 4.47 4.45 3.86 3.88 3.78 3.55 3.69 3.81 3.75 3.98 4.22 3.98
Qwen3-8B 4.46 4.51 4.50 3.99 3.90 3.78 3.75 3.87 3.95 3.80 3.97 4.15 4.05
Llama-3.1-8B-Instruct 4.22 4.27 4.28 3.69 3.65 3.65 3.48 3.61 3.74 3.84 3.95 4.22 3.88
Table 4: Performance of models at each stage of the Emotion Conversation task under the Base setting. The entire conversation is divided into four stages: i) Reception and Inquiry, ii) Diagnostic, iii) Consultation, and iv) Consolidation and Conclusion. Each stage includes 3 checkpoints, denoted as X-Y, where X indicates the stage number and Y indicates the checkpoint index.
In the Emotion Classification, Emotion Detection, Emotion QA, and Emotion Expression, we employ GPT-4o as the CoEM-Sage, while Deepeek-V3 is used for the Emotion Conversation-4 and Emotion Summary in the same role. For the retrieval and ranking components across both the RAG and CoEM settings, we adopt bge-m3 (Chen et al. 2024a) as the CoEM-Rank. The generation models listed in Table 3 are used as the CoEM-Core. Full configuration details for both the RAG and CoEM frameworks are in Appendix E.
@@ 5.2 Results on LongEmotion
The overall experimental results can be seen in Table 3. We evaluate the performance of each model on all tasks under the Base, RAG, and CoEM settings. As the first three stages of the dialogue are relatively brief, RAG and CoEM are only applied in the fourth stage of the Emotion Conversation task. The performance of models in all stages under the Base setting can be seen in Table 4.
By analyzing the experimental results, we can observe the following: i) While GPT-4o and DeepSeek-V3 generally exhibit stronger Emotional Intelligence, Llama-3.1-8B-Instruct and Qwen3-8B significantly outperform GPT-4o and GPT-4o-mini in the Emotion Conversation-4 task. This is further supported by our experimental results across all stages in the Base setting, as shown in Table 4. ii) In the Emotion Classification and Emotion Detection tasks, which heavily test the models’ reasoning and classification abilities, we maximize the potential of the models through the use of CoEM. iii) In contrast, in the Emotion QA and Emotion Summary tasks, which are strongly context-based, the model’s score largely depends on the alignment between the model’s response and the original text. Therefore, injecting external knowledge may introduce harmful noise into the context, leading to a drop in the score. iv) In the Emotion Expression task, we use GPT-4o as the CoEM-Sage to enrich the model’s expression. Compared to the results of RAG and CoEM, the score of GPT-4o-mini drops, while the scores of the other four models improve. This indicates that the ability of the CoEM-Sage greatly influences the performance of the tested models. Our ablation study on the CoEM-Sage models for Emotion Detection and Emotion Summary further supports this conclusion, as shown in Figure 5.
@@
To explore models’ ability in emotion recognition across different context lengths, we evaluate their performance on the Emotion Classification under the Base setting, as shown in Figure 6. It can be observed that GPT-4o demonstrates the overall best performance, while DeepSeek-V3 shows the highest stability. In the longest range of 24k-27k, Llama-3.1-8B-Instruct experiences a significant drop in performance, reflecting its limitations in handling long contexts.
Figure 6: Model accuracy by context length range on Emotion Classification.
We also conduct ablation experiments on RAG with different chunk sizes and retrieval quantities, as shown in Figure 7. From the image, it can be seen that GPT-4o-mini achieved the best performance in the 128 tokens/chunk setting with 8 retrieved chunks. Furthermore, although increasing the chunk size or retrieved count allows the model to acquire more information, it also introduces more noise, which can harm the model’s performance. Therefore, selectively incorporating useful information and discarding irrelevant information is crucial to improving RAG performance.
@@
In constructing the Emotion Conversation dataset, we employ the CPsyCoun two-stage data‑generation framework and enhance our prompt design. To illustrate the richness of emotional features in our synthetic data, we conduct comparative quality experiments against plain prompts. The experimental results are detailed in the Appendix D.
5.3 Comparison of GPT series models
@@
In the Emotion Summary task, GPT-4o-mini and GPT-4o directly analyzed various features of the case, whereas GPT-5 structured its analysis based on psychological theories, resulting in a higher score, as shown in Figure 11.
@@
From the tasks above, we can conclude that GPT-4o-mini behaves more like a human, with richer emotional features, but its application of psychological theory is somewhat lacking. On the other hand, GPT-5 has a better understanding of psychological theories, but the output is too rigid and mechanical, which might lead to a less empathetic user experience in practice. GPT-4o strikes a more balanced approach between theoretical understanding and emotional features.
6 Conclusion
In this work, we introduce LongEmotion, a benchmark for measuring models’ Emotional Intelligence in long-context scenarios. LongEmotion comprises six tasks that comprehensively challenge models across multiple dimensions—emotion recognition, emotional support, emotional expression, emotional knowledge, and more. Beyond constructing the dataset, we also build Retrieval-Augmented Generation (RAG) and Collaborative Emotional Modeling (CoEM) frameworks for each task, achieving improvements on the vast majority of them. We conduct exhaustive experiments on the LongEmotion dataset under Base, RAG, and CoEM settings, analyzing models’ Emotional Intelligence from perspectives such as emotion enhancement, long-text performance, and expressive capability. Additionally, we integrate rigorous manual annotations into our synthetic data creation pipeline to ensure high data quality.
7 Future Work
In future work, we will gradually supplement the performance of new open-source and closed-source models under the base setting, and open-source all data, code, and evaluation results.
Appendix A Qualifications of Annotators
Our annotation team consists of psychology researchers and computer science researchers. In the psychology research team, there is a postdoctoral fellow expert specializing in psychology and seven Master’s students majoring in the same field. The theoretical foundation of our dataset and metrics involves deep participation from the psychology team. Under the guidance of the expert, the seven psychology Master’s students carry out the annotation work. In the computer science research team, there are three Master’s students and one PhD student majoring in computer science. Their main responsibility is to modify, adjust, and organize the data annotated by the psychology team according to the characteristics of the tasks.
Appendix B Inter-annotator Agreement
We use inter-annotator agreement to measure the consistency among human annotators. Specifically, our annotators independently re‑annotate the same set of 20 Emotion Conversation examples—yielding a total of 240 metric‑level judgments. We calculate inter-annotator agreement using Fleiss’ Kappa coefficient, with results presented in Table 5.
Metric Result
Establishing the Therapeutic Alliance -0.064
Emotional Acceptance and Exploration Guidance 0.037
Systematic Assessment -0.156
Recognizing Surface-Level Reaction Patterns 0.045
Deep Needs Exploration -0.005
Pattern Interconnection Analysis -0.011
Adaptive Cognitive Restructuring 0.004
Emotional Acceptance and Transformation -0.057
Value-Oriented Integration -0.055
Consolidating Change Outcomes and Growth Narrative -0.065
Meaning Integration and Future Guidance -0.111
Autonomy and Resource Internalization -0.022
Table 5: Fleiss’ kappa coefficient for inter-annotator agreement in Emotion Conversation.
Appendix C Case Study
In this section, we conduct a concrete analysis of how the information retrieved by the RAG and CoEM methods affects model performance. We collect all of the information retrieved by RAG and CoEM for every task in Figure 12. In models’ final generation prompts, the Base setting includes none of the information; the RAG setting includes only the Content information; and the CoEM setting includes both the Content and Summary information.

### v2

@@ 4 Collaborative Emotional Modeling
Figure 5 illustrates the pipeline of CoEM. To address EI tasks involving long contexts, we propose a hybrid retrieval-generation architecture that combines Retrieval-Augmented Generation (RAG) with modular multi-agent collaboration. For the parameter settings and application details, please refer to Appendix C. For the case analysis of RAG and CoEM, please refer to Appendix B.2. The framework consists of five key stages:
Figure 5: The pipeline of Collaborative Emotional Modeling (CoEM).
@@
The context is segmented into token-length-constrained chunks, whereas in Emotion Detection, each segment is considered as an individual chunk. We set different chunk sizes based on the characteristics of each task. We demonstrate the parameter settings in Appendix C.
@@
A retrieval agent, implemented as CoEM-Rank, evaluates the relevance of each chunk to the query using dense semantic similarity, with relevance scores computed based on cosine similarity. Top-ranked chunks are passed forward for enhancement. By ranking the original context chunks, the factual relevance of the retrieved information is ensured.
@@
A reasoning agent called CoEM-Sage, functioning as a knowledge assistant, enriches the selected chunks by incorporating external knowledge or latent emotional signals through our task-specific prompts. Specifically, in Emotional Recognition tasks, CoEM-Sage identifies subtle emotional cues; in Knowledge Application tasks, it provides summaries based on psychological knowledge; and in Empathetic Generation tasks, it enhances CoEM-Core’s empathy and expression through emotional analysis. These signals, derived from psychological theories or curated priors, are incorporated into the original chunks without task-specific leakage.
@@
The enriched chunks, now augmented with emotional features, are then re-evaluated by CoEM-Rank for their semantic relevance to the query, measured by cosine similarity. This final ranking ensures that the selected context is not only factually grounded but also affectively coherent. By ranking the enriched chunks, the emotional relevance of the retrieved information is ensured, as these chunks contain not only the original text but also external emotional information.
@@
The selected and enriched chunks, along with the context and prompt, is fed into a generation model denoted as CoEM-Core. This model (e.g., a long-context LLM or an instruction-tuned model) produces the final task-specific output, whether it be classification, summarization, or dialogue generation.
This modular approach encourages interpretability, emotional awareness, and task robustness. The CoEM setting encompasses all five stages, while the RAG setting only comprises Chunking, one-time Ranking, and Emotional Ensemble Generation. We conduct an empirical case study of the framework, which can be found in Appendix B.2.
@@ 5.1 Experiment Setup
In our experiments, for closed-source models, we choose GPT-4o-mini (OpenAI, 2024a) and GPT-4o, while for open-source models, we select DeepSeek-V3 (DeepSeek-AI, 2024), Llama3.1-8B-Instruct (Grattafiori et al., 2024), and Qwen3-8B (Team, 2025). For tasks employing automatic evaluation, we adopt GPT-4o as the evaluator. Under the base setting, we compare a broader range of advanced open-source and closed-source models. For comparison, we have the performance of GPT-5 (OpenAI, 2025), Qwen3-14B and Qwen3-32B under the Base setting.
To accelerate inference, we use vllm library (Kwon et al., 2023) as the inference engine and set temperature=0.8 and top_p=0.9 for all open-source models. For Qwen3 series models, we enable its thinking capabilities and manually remove the reasoning process between <think> and </think> to keep the answers concise. All experiments are conducted using NVIDIA A800 80G GPUs, with open-source models under 14B parameters running on a single GPU and the 32B models utilizing two GPUs.
Method Model Recognition Knowledge Generation Overall
EC ED QA ES MC-4 $E$ Avg
Base GPT-4o-mini 37.00 16.42 48.61 4.54 3.75 86.77 59.10
GPT-4o 50.09 19.12 50.12 4.60 3.77 81.03 61.29
DeepSeek-V3 56.50 24.51 45.53 4.62 3.99 81.75 63.42
Qwen3-8B 48.00 18.14 44.75 4.51 3.97 73.40 58.98
Llama3.1-8B-Instruct 39.34 9.80 44.56 4.29 4.00 75.61 55.85
(Extended Comparison Models)
GPT-5 73.75 22.79 43.22 4.42 4.67 86.77 68.06
Qwen3-14B 50.00 20.83 46.35 4.55 3.95 84.49 61.95
Qwen3-32B 58.25 20.59 43.11 4.53 4.17 84.81 63.46
RAG GPT-4o-mini 51.67 21.57 50.72 4.53 3.78 80.41 61.76 $\uparrow$2.66
GPT-4o 61.34 22.55 51.81 4.52 3.80 79.49 63.60 $\uparrow$2.31
DeepSeek-V3 62.59 23.53 50.44 4.63 4.34 81.83 66.30 $\uparrow$2.88
Qwen3-8B 41.59 19.12 44.34 4.54 4.14 73.28 58.65 $\downarrow$0.33
Llama3.1-8B-Instruct 44.00 11.27 43.21 4.26 3.94 75.16 56.27 $\uparrow$0.42
CoEM GPT-4o-mini 59.50 20.59 49.12 4.52 3.77 80.38 62.57 $\uparrow$3.47
GPT-4o 61.42 25.00 51.07 4.53 3.81 80.41 64.12 $\uparrow$2.83
DeepSeek-V3 64.17 23.04 50.39 4.65 4.34 82.83 66.70 $\uparrow$3.28
Qwen3-8B 62.92 18.14 51.11 4.55 4.14 73.59 63.26 $\uparrow$4.28
Llama3.1-8B-Instruct 55.09 11.27 44.79 4.17 4.00 75.71 58.38 $\uparrow$2.53
Table 2: Experiment result across Base, RAG and CoEM. MC-4 represents the fourth stage of Emotion Conversation. By aligning the MC-4 and ES scores with the 100-point scale, the overall score is computed as (EC + ED + $E$ + QA + MC-4×20 + ES×20)/6, where the numbers to the right indicate the score change relative to the Base setting.
In the EC (Emobench as needle), ED, and EE, we employ GPT-4o as the CoEM-Sage, while DeepSeek-V3 is used for the EC (Finentity as needle), QA, MC and ES in the same role. For the retrieval and ranking components across both the RAG and CoEM settings, we adopt bge-m3 (Chen et al., 2024a) as the CoEM-Rank. The generation models listed in Table 2 are used as the CoEM-Core. Configuration details for both the RAG and CoEM frameworks are in Appendix C.
@@ 5.2 Results on LongEmotion
The overall experimental results can be seen in Table 2. We evaluate the performance of each model on all tasks under the Base, RAG, and CoEM settings. As the first three stages of the dialogue are relatively brief, RAG and CoEM are only applied in the fourth stage of the Emotion Conversation.
Overall Analysis of Experimental Results.
As shown in Table 2, DeepSeek-V3 and GPT models exhibit generally strong EI capabilities, achieving stable performance gains even with vanilla RAG. In contrast, Qwen3-8B and Llama-3.1-8B-Instruct perform less effectively under the RAG setting, suggesting that some models struggle to effectively integrate retrieved chunks within long-context reasoning. This limitation can be mitigated by CoEM, which enhances contextual alignment and emotional reasoning through multi-agent collaboration.
Ablation Experiments.
To evaluate the effectiveness of RAG-based methods in enhancing EI, we integrate Self-RAG and Search-o1 into LongEmotion using Qwen3-8B as the base model. In the Self-RAG setting, retrieved chunks are rescored by Self-RAG-7B for relevance, with irrelevant ones filtered out before concatenation with the prompt. The additional use of Self-RAG-7B outputs in the ES task further improves performance, showing that selectively enriching retrieved information benefits emotional intelligence. In the Search-o1 setting, Qwen3-8B autonomously generates queries and retrieves relevant chunks via Bge-m3 within five search turns. The observed performance drop indicates that small-scale models struggle with autonomous search-based reasoning in emotional tasks. Results are reported in Table 3.
Method EC ED QA ES MC-4
RAG 41.59 19.12 44.34 4.51 4.14
Self-RAG 44.00 16.18 44.02 4.57 4.15
Search-o1 45.25 16.18 45.12 4.50 3.72
CoEM 62.92 18.14 51.11 4.55 4.14
Table 3: Ablation experiment results on methods.
To investigate how the reasoning processes of models affect their Emotional Intelligence in long-context scenarios, we perform ablation studies on the Qwen3 model series using two emotion recognition tasks—Emotion Classification (Emobench as Needle) and Emotion Detection—along with one empathetic generation task, Emotion Expression, under the Base setting. By analyzing Table 4, we can observe that through thinking, Qwen3-8B achieve the most significant improvement, while the improvement of Qwen3-14B is not substantial.
Task Qwen3-8B Qwen3-14B Qwen3-32B
think w/o think w/o think w/o
EC-E 38.50 28.67 31.00 30.75 48.00 37.50
ED 18.14 12.01 20.83 20.83 20.59 20.10
$E$ 73.40 70.32 84.49 83.13 84.81 84.02
Table 4: Ablation experiments of the thinking process in the Qwen3 series models.
Furthermore, to examine how the capability of CoEM-Sage affects the overall framework, we perform ablation experiments on the MC-4 task. As shown in Table 2, DeepSeek-V3 outperforms GPT-4o under the base setting. Consistently, when used as the CoEM-Sage, DeepSeek-V3 also drives higher performance than GPT-4o, as can be seen in Figure 6. These results further demonstrate the soundness and scalability of CoEM.
Figure 6: Impact of CoEM-Sage models on MC-4.
@@
To explore models’ ability in emotion recognition across different context lengths, we evaluate their performance on the Emotion Classification (Finentity as Needle) under Base setting, as shown in Figure 7. DeepSeek-V3 and Qwen3-8B exhibit both high stability and strong overall performance, whereas GPT-based models show weaker robustness in long-context settings, in some cases even performing below Llama-3.1-8B-Instruct.
Figure 7: Model accuracy by context length on EC.
We further conduct ablation experiments on RAG with varying chunk sizes and retrieval counts, as shown in Figure 8. GPT-4o-mini performs best with 128-token chunks and eight retrieved segments, while larger settings introduce noise and reduce overall performance.
@@
Case Study.
(i) First, we qualitatively compare the GPT model series across all tasks under the Base setting, revealing that GPT-5 is theoretically stronger but more mechanical and prone to hallucination, GPT-4o-mini exhibits more human-like behavior yet lacks theoretical grounding, while GPT-4o achieves a balanced trade-off. (ii) Furthermore, we visualize the CoEM framework and empirically analyze its influence on emotional information. (iii) Finally, we analyze the advantages of the LongEmotion dataset in advancing Emotional Intelligence. For complete details of case study, please refer to Appendix B.
6 Conclusion
In this work, we introduce LongEmotion, a benchmark for measuring models’ Emotional Intelligence in long-context scenarios. LongEmotion comprises six tasks that comprehensively challenge models across emotion recognition, knowledge application and empathetic generation. Beyond constructing the dataset, we also build Retrieval-Augmented Generation (RAG) and Collaborative Emotional Modeling (CoEM) frameworks for each task, achieving improvements on the vast majority of them. We conduct exhaustive experiments and a detailed case study to analyze models’ EI in long-context scenarios.
7 Limitations
In this work, we propose LongEmotion, a benchmark for evaluating the emotional intelligence of LLMs in long-context inference. However, all the datasets in our benchmark are based solely on the text modality and are restricted to the psychological and emotional domains. Similarly, the proposed CoEM framework focuses only on textual inputs and does not extend to other modalities such as vision or audio. In addition, our dataset includes only English texts and does not cover other languages. It remains uncertain whether the same level of quality can be preserved when the data are translated into other languages.
8 Ethical Considerations
Data Privacy
In this work, all the datasets we adopt are formally published in academic venues and comply with data privacy and ethical protection standards. Through data augmentation and manual inspection, we ensure that no ethical risks are introduced. In addition, all annotators involved in our dataset construction possess academic backgrounds in computer science or psychology, ensuring the reliability of the data annotation process. We adhere to the intended use and license terms of all source datasets. The datasets in LongEmotion are intended solely for academic research and will not be used for any other purposes.
Potential Risks
All models evaluated in our experiments, including both open-source and closed-source ones, are officially released models, which helps ensure that no harmful or unsafe content is generated. In addition, all prompts used in our evaluation are fully disclosed in this paper, which can be seen in Appendix F, and these prompts are carefully designed to ensure a high level of safety.
Appendix A Qualifications of Annotators
Our annotation team consists of psychology researchers and computer science researchers. In the psychology research team, there is a postdoctoral fellow expert specializing in psychology and seven Master’s students majoring in the same field. The theoretical foundation of our dataset and metrics involves deep participation from the psychology team. Under the guidance of the expert, the seven psychology Master’s students carry out the annotation work. In the computer science research team, there are three Master’s students and one PhD student majoring in computer science. Their main responsibility is to modify, adjust, and organize the data annotated by the psychology team according to the characteristics of the tasks. All student annotators and researchers involved in the annotation and data processing work receive reasonable financial compensation for their time and effort, commensurate with local standards and the complexity of the tasks.
Appendix B Case Study
B.1 Comparison of GPT series models
@@
In the Emotion Summary task, both GPT-4o-mini and GPT-4o directly analyze various features of the case, while GPT-5 structures its analysis based on psychological theories. However, GPT-5 exhibits hallucinations, often adding non-existent facts. For instance, in Figure 12, the term "slapping" is highlighted in red, but the source data never mentions such an action.
@@
From the tasks above, we can conclude that GPT-4o-mini behaves more like a human, with richer emotional features, but its application of psychological theory is somewhat lacking. On the other hand, GPT-5 has a better understanding of psychological theories, but the output is too rigid and mechanical, which might lead to a less empathetic user experience in practice. Additionally, GPT-5 tends to exhibit hallucinations, often adding non-existent facts. GPT-4o strikes a more balanced approach between theoretical understanding and emotional features.
B.2 Case Analysis of RAG and CoEM
We conduct a concrete analysis of how the information retrieved by the RAG and CoEM methods affects model performance. In models’ final generation prompts, the Base setting includes none of the information; the RAG setting includes only the Chunk information; and the CoEM setting includes both the Chunk and Summary information.